OpenStudy (samsan9):

find the center and the radius of x^2+y^2+6x-16y+48=0

OpenStudy (zzr0ck3r):

Do you know how to complete the square?

OpenStudy (samsan9):

um is that when you factor it perfectly?

OpenStudy (zzr0ck3r):

\(x^2+y^2+6x-16y+48=0\\x^2+6x+y^2-16y+48=0\\(x+3)^2-9+y^2-16y+48=0\)

OpenStudy (zzr0ck3r):

does this make sense?

OpenStudy (zzr0ck3r):

\(x^2+bx = (x+\frac{b}{2})^2-(\frac{b}{2})^2\)

OpenStudy (zzr0ck3r):

so \(x^2+6x = (x+3)^2-9\)

OpenStudy (zzr0ck3r):

similarly \(y^2-16y = (y-8)^2-64\)

OpenStudy (samsan9):

yes so far it is

OpenStudy (zzr0ck3r):

so \(x^2+6x+y^2-16y+48=0\\(x+3)^2-9+(y-8)^2-64+48=0\\(x+3)^2+(y-8)^2-25=0\\(x-(-3))^2+(y-8)^2=25\\(x-(-3))^2+(y-8)^2=5^2\)

OpenStudy (zzr0ck3r):

the equation of a circle with center \((a,b)\) and radius \(r\) is \((x-a)^2+(y-b)^2=r^2\)

OpenStudy (samsan9):

so is the center is (-3,8) ?

OpenStudy (zzr0ck3r):

correct

OpenStudy (samsan9):

and the radius is just 5?

OpenStudy (zzr0ck3r):

correct